# A Nuclear Interferometer for Ultra-Light Dark Matter Detection

### v1

@@ V Conclusions
With an extraordinarily broad landscape of viable candidates which give rise to a diverse range of phenomenological signatures, new ideas for the detection of DM are required in order to maximise experimental coverage of the vast available parameter space. In this work we propose two possible realisations of a single-photon interferometry experiment based on the nuclear clock transition in 229Th as a probe of ULDM with scalar couplings to the SM, making use of single-ions and clouds of neutral atoms respectively. Although working with either source type presents a unique set of experimental challenges, we find that the enhanced sensitivity of the nuclear clock transition to the variation of fundamental constants offers the means to access large regions of uncharted parameter space over a wide range of frequencies.
@@ Acknowledgements.
This work is supported by the CERN Quantum Technology Initiative. We are incredibly grateful to Leondardo Badurina and Thomas Hird for discussions regarding the challenges of using sources with short excited state lifetime in atom interferometry. We also thank Dennis Schlippert, Marianna Safranova and Gilad Perez for invaluable feedback on early versions of this manuscript, Ekkehard Peik and Michael Doser for insight on the nuclear clock and John Carlton for discussions on the AION experiment. We thank Nathaniel Sherrill and Abhishek Banerjee for details on the analyses of atomic clock comparison experiments. HB acknowledges partial support from the STFC HEP Theory Consolidated grants ST/T000694/1 and ST/X000664/1 and thanks other members of the Cambridge Pheno Working Group for useful discussions. EF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967.
@@ Appendix B $\pi$-pulse efficiencies
Appendix C Photo-ionisation vs. photo-excitation of neutral 229Th atoms
The first ionisation potential of neutral 229Th is 6.3 eV - lower than the 8.3 eV nuclear isomeric excitation energy. With the interferometry schemes proposed in this work reliant on the successful completion of a precise sequence of coherent, laser-induced nuclear excitations and de-excitations, it is necessary to explore the extent to which photo-ionisation would compete with these processes, depleting the number of atoms at the point of detection and reducing the sensitivity. By comparing estimations of the photo-ionisation and nuclear photo-excitation cross-sections of atomic 229Th, we offer a rough assessment of this here before exploring a possible means of mitigation.
From the ground electronic configuration, a laser resonant with the isomeric transition is sufficiently energetic to ionise the atom via removal of electrons from either the 6d or 7s electronic shells. At leading order, these processes proceed via allowed (E1) electric dipole transitions with a rate that is governed by the scalar product of the laser polarisation with the effective electric dipole moment of the transition (i.e. the promotion of the bound electron to the continuum) in question [128]. In an ensemble of unpolarised atoms, the electric dipole transition moment of a given process is randomly orientated in different atoms. In this instance, the photo-ionisation rate follows from the average over all possible orientations of the transition dipole moment relative to the laser polarisation. If the sample is polarised however, these moments align in a direction fixed by the electromagnetic field used to polarise the ensemble. As demonstrated in Ref. [129], the measured rate of that particular photo-ionisation process then drops to zero if the laser polarisation is chosen to be orthogonal to the direction of the transition moment.
The cross-sections for E1 photo-ionisation of the 6d and 7s electrons are computed numerically in Ref. [130] (assuming the atoms to be unpolarised) using the flexible atomic code (FAC) based upon the relativistic configuration interaction method [131]. At an energy of $\sim$ 8 eV these cross-sections are of order 10-17 cm2 [130].
In the low saturation limit of moderate laser intensities and a laser bandwidth that exceeds the natural linewidth of the transition, the resonance nuclear photo-excitation cross-section is given by [127]
where $\Gamma_{\gamma}\sim 10^{-9}\hskip 5.0pt\Gamma_{a}$ is the radiative decay rate and $\lambda_{0}$ is the transition wavelength. Setting $\Gamma_{l}$ to $\Gamma_{a}$, its minimum value consistent with the regime for which this equation is valid, gives $\sigma_{\textnormal{ex.}}\sim 10^{-20}$ cm2, approximately three orders of magnitude below the photo-ionisation rate. Whilst, as stated in Appendix B, our sensitivity forecasts assume the laser bandwidth to be narrow relative to the natural linewidth to reduce atom loss from spontaneous decay, here we are just interested in an order of magnitude estimate of the nuclear excitation rate in order to assess the extent to which photo-ionisation is likely to be a problem. Given that the nuclear excitation rate should be greatest in the low saturation limit, using this result provides a conservative assessment of the potential impact of photo-ionisation on the interferometry sequence and should thus be sufficient here.
As such, we conclude that performing interferometry with unpolarised clouds of 229Th atoms is unlikely to be viable due to a significant depletion of the number of atoms remaining coherently in the cloud at detection. To perform nuclear interferometry with neutral 229Th atoms, we suggest that it is therefore necessary for them to be polarised such that the transition electric dipole moments of different atoms become aligned. By then making use of a linearly polarised laser with a polarisation carefully chosen to be orthogonal to the transition dipole moments of the two E1 photo-ionisation processes, it could therefore be possible to forbid photo-ionisation at leading order. Whilst photo-ionisation could still proceed via higher order (e.g. first forbidden magnetic dipole M1) processes, these rates are, according to the computations in Ref. [130], approximately 8 orders of magnitude smaller than the E1 processes and are thus negligible with respect to estimations of the nuclear photo-excitation rate. The forecasts presented in this work neglect photo-ionisation, assuming it has been suppressed to negligible levels by successful adoption of such a practice.
Appendix D AION Prospective Sensitivity

### v2

@@ V Conclusions
With an extraordinarily broad landscape of viable candidates which give rise to a diverse range of phenomenological signatures, new ideas for the detection of DM are required in order to maximise experimental coverage of the vast available parameter space. In this work we propose two possible realisations of a single-photon interferometry experiment based on the nuclear clock transition in 229Th as a probe of ULDM with scalar couplings to the SM, making use of single-ions and clouds of neutral atoms respectively. Although working with either source type presents a unique set of experimental challenges, we find that the enhanced sensitivity of the nuclear clock transition to the variation of fundamental constants offers the means to access large regions of uncharted parameter space over a wide range of frequencies providing experimental challenges can be met in the future.
@@ Acknowledgements.
This work is supported by the CERN Quantum Technology Initiative. We are incredibly grateful to Leondardo Badurina and Thomas Hird for discussions regarding the challenges of using sources with short excited state lifetime in atom interferometry and to Sebastian Ellis for raising the issue of magnetic field noise after submission of V1, reflected in the extended discussion in this V2. We also thank Dennis Schlippert, Marianna Safranova, Masha Baryakhtar, Gilad Perez and Ekkehard Peik for invaluable feedback on early versions of this manuscript, Michael Doser for discussions on the nuclear clock and John Carlton for insight into the AION experiment and gravitational gradient noise. We thank Jeremiah Mitchell for discussions regarding the treatment of noise sources and Nathaniel Sherrill and Abhishek Banerjee for details on the analyses of atomic clock comparison experiments. HB acknowledges partial support from the STFC HEP Theory Consolidated grants ST/T000694/1 and ST/X000664/1 and thanks other members of the Cambridge Pheno Working Group for useful discussions. EF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2123 QuantumFrontiers – 390837967, and via the SFB/CRC 1227 (DQ-mat) – Project-ID 274200144 – of the DFG.
@@ Appendix B $\pi$-pulse efficiencies
Appendix C AION Prospective Sensitivity